# Spectuner-D1: Spectral Line Fitting of Interstellar Molecules Using Deep Reinforcement Learning

### v1

@@ Abstract
Spectral lines from interstellar molecules provide crucial insights into the physical and chemical conditions of the interstellar medium. Traditional spectral line analysis relies heavily on manual intervention, which becomes impractical when handling the massive datasets produced by modern facilities like ALMA. To address this challenge, we introduce a novel deep reinforcement learning framework to automate spectral line fitting. Using observational data from ALMA, we train a neural network that maps both molecular spectroscopic data and observed spectra to physical parameters such as excitation temperature and column density. The neural network predictions can serve as initial estimates and be further refined using a local optimizer. Our method achieves consistent fitting results compared to global optimization with multiple runs, while reducing the number of forward modeling runs by an order of magnitude. We apply our method to pixel-level fitting for an observation of the G327.3-0.6 hot core and validate our results using xclass. We perform the fitting for CH3OH, CH3OCHO, CH3OCH3, C2H5CN, and C2H3CN. For a 100 $\times$ 100 region covering 5 GHz bandwidth, the fitting process requires 4.9 to 41.9 minutes using a desktop with 16 cores and one consumer-grade GPU card.
@@
facilities: ALMA, SMA, IRAM:30msoftware: astropy (2013A&A...558A..33A; 2018AJ....156..123A; 2022ApJ...935..167A), casa (2022PASP..134k4501C), matplotlib (Hunter:2007), numpy (harris2020array), pandas (mckinney-proc-scipy-2010; reback2020pandas), scipy (2020SciPy-NMeth), statcont (2018A&A...609A.101S), xclass (2017A&A...598A...7M).
1 Introduction
@@
Spectral lines from interstellar species provide a crucial probe of the physical and chemical conditions in the interstellar medium (ISM). The analysis of spectral lines involves several key steps: querying transitions from spectroscopic databases, matching these transitions to observed spectral lines, and determining a set of physical parameters that best reproduce the observed spectrum (e.g. 1999ApJ...517..209G). Traditionally, spectral line analysis has been a labor-intensive process, requiring significant expertise and time to manually match observed features with known transitions and iteratively adjust parameters to achieve satisfactory fits. Although several analysis tools have been developed over the past decade, e.g. weeds (2011A&A...526A..47M), cassis (2015sf2a.conf..313V), xclass (2017A&A...598A...7M), madcuba (2019A&A...631A.159M), and pyspeckit (2022AJ....163..291G), manual intervention remains common. As the volume and complexity of observational data continue to increase due to advanced observational facilities like the Atacama Large Millimeter/submillimeter Array (ALMA) (2020arXiv200111076C), there is a growing need for automated and efficient methods to perform this analysis, especially in line-rich regions such as hot cores and hot corinos (see 2020ARA&A..58..727J, for a review).
@@
Recent advances in machine learning offer a promising opportunity to transform spectral analysis workflows. For instance, machine learning algorithms have been used to establish the relation between spectral line properties to the underlying physical parameters (2024A&A...691A.109E; 2025A&A...702A..71G; 2025A&A...698A.286M). 2025arXiv251009119K trained a convolutional neural network for identification of complex organic molecules (COMs) using their spectral line features. In addition, machine learning techniques were applied to explore the chemical inventories of molecular clouds (2021ApJ...917L...6L; 2025ApJ...985..123T) and to investigate the interstellar chemical reaction networks (2022A&A... [^6] A..45V; 2023MNRAS.526..404H).
@@
In this paper, we present a deep learning framework to automate and enhance the spectral line fitting process. Specifically, we aim to train a neural network that can infer the physical parameters of molecule emission such as excitation temperature and column density from the observed spectrum and the molecular transition properties within the observed spectral windows. We choose deep learning over traditional machine learning methods such as random forests (breiman2001random) because it scales more effectively with large datasets and is better suited to handling the massive data volumes produced by ALMA and other modern observatories. The inferred physical parameters of the neural network are based on a one-dimensional radiative transfer model that assumes local thermodynamic equilibrium (LTE). While this assumption may introduce bias (e.g. 2024A&A...686A.255R), the simplicity of the LTE model enables efficient computation of model spectra, which significantly reduces the time required for training and testing and makes the project feasible.
@@
This work employs deep reinforcement learning (DRL) (e.g. mnih2015human; schulman2017proximal) to train the neural network. To improve the generalization of the neural network, we seek to train it using real ALMA data. However, vast amounts of labeled data are unavailable for this task. This challenge could be addressed using DRL. Unlike supervised learning, DRL does not require labeled data and instead discovers optimal results during training. In our approach, training proceeds by interacting with a reward model and receiving feedback through reward signals that reflect fit quality. The reward model uses the physical parameters inferred by the neural network to compute the model spectra under LTE conditions and estimates the fit quality by comparing them to the observed spectrum.
@@
Our neural network-based approach can be efficiently applied to pixel-by-pixel fitting of spectral line cubes, which is a challenging task. When using local optimization algorithms, providing good initial parameter estimates for each pixel is non-trivial, and poor initial guesses can significantly affect the the fitting results. For instance, to improve the fitting, 2024ApJ...965...14E proposed using the results from previously fitted neighboring pixels to initialize the parameters for subsequent fits. On the other hand, while global optimization methods could lead to more robust fitting results, they are time consuming and require large number of forward modeling runs. In contrast, our neural network can provide good initial guesses for all pixels, leading to efficient and robust pixel-by-pixel fitting.
This paper is organized as follows. Section 2 introduces our methodology, including the description of the training data, along with the training and evaluation methods. Section 3 presents the training results. In Section 4, we apply our method to pixel-level spectral fitting. We then discuss our methodology in Section 5. Finally, this work is summarized in Section 6.
@@
Figure 1: Architecture of the proposed neural network introduced in Section 2.2. Here, FC stands for the fully connected layer, MLP stands for the multi-layer perceptron, and SwiGLU stands for the SwiGLU activation function (2020arXiv200205202S).
2 Methodology
2.1 Network inputs
@@
Given an observed spectrum with multiple spectral windows and a molecule identifier, we first query the molecular spectroscopic database to find all transitions of the molecule within those frequency ranges, resulting in a $N_{\text{trans}}\times F_{\text{trans}}$ array, denoted as $\textbf{d}_{\text{trans}}$, where $N_{\text{trans}}$ is the number of transitions and $F_{\text{trans}}$ is the number of features of each transition. For example, if the input spectral windows span 97 - 99 GHz and 100 - 102 GHz, and the target molecule is CH3OH, v=0, we query the molecular spectroscopic database for all transitions of CH3OH, v=0 within these frequency ranges, and may find 25 transitions; in this case $N_{\text{trans}}=25$. Data are grouped into batches during training. A single batch may contain multiple species, e.g. CO, v=0, CH3OH, v=0, and H2CO, v=0, and the corresponding spectral windows may also be different. Therefore, the number of transitions $N_{\text{trans}}$ varies across these input data. This situation is common in NLP, where the number of transitions is analogous to sentence length, and can be effectively handled using transformer-based neural networks (NIPS2017_3f5ee243). We adopt this network architecture in our work, as described in Section 2.2.
We include the following features for each transition: the energy of the lower state $E_{\text{l}}$, the energy of the upper state $E_{\text{u}}$, the upper state degeneracy $g_{\text{u}}$, the Einstein coefficient $A_{\text{ul}}$, and the partition function Q at different temperatures. The molecular spectroscopic data are from the Virtual Atomic and Molecular Data Center (VAMDC) (2016JMoSp.327...95E), which contains entries from the Cologne Database for Molecular Spectroscopy (CDMS) (2001A&A...370L..49M; 2005JMoSt.742..215M) and the Jet Propulsion Laboratory (JPL) database (1998JQSRT..60..883P).
@@
The third feature is derived using the approach described in 2025ApJS..277...21Q. Specifically, we use the find_peaks function implemented by scipy, which identifies all local maxima in a 1D input array based on user-specified criteria. The most relevant criterion of the algorithm is the prominence threshold. Prominence is a property of a peak, defined as the vertical distance from the peak height to the lowest contour line that does not enclose any higher peak. The algorithm ignores any peak with a prominence below the specified threshold. In this work, we set the prominence threshold to four times the root mean square (RMS) noise of the input spectrum. The method to estimate the noise levels is described in Section 2.4.1. In addition, the peak finding algorithm assumes that the input spectrum has a flat baseline. We have checked that this assumption is valid for the vast majority of the observational data used in this study.
@@
Table 1: Summary of the fitting parameters in our spectral line model described in Section 2.3.1.
@@ 2.2 Network architecture
Our network architecture, which consists of an encoder and a decoder, is illustrated in Figure 1. The neural network only uses basic modules in deep learning, including the fully connected (FC) layer, the multilayer perceptron (MLP), the patching layer in the vision transformer (ViT) (dosovitskiy2021an), the SwiGLU activation function (2020arXiv200205202S), the multi-head attention module (NIPS2017_3f5ee243) and the transformer layer (NIPS2017_3f5ee243). Readers may also refer to Goodfellow-et-al-2016 for the relevant concepts of deep learning.
@@
The encoder of our neural network converts the input variable-length sequences of molecular spectroscopic and observational data into dense vectors. This process is analogous to models in natural language processing (NLP) that learn vector embeddings of sentences and paragraphs, such as BERT (devlin-etal-2019-bert). A transformer-based architecture is particularly suitable for this purpose, and the so-called [CLS] token (devlin-etal-2019-bert) can be used to represent aggregate sequence-level information. In astronomy, transformer-based neural networks have been applied to learning the representations of galaxy images and spectra (e.g. 2024MNRAS.531.4990P; 2025AJ....170...28R).
@@
For the sequence of the observed spectrum, we first split the sub-window spectrum into several patches using the patching layer in ViT (dosovitskiy2021an). The same patching operation is applied across the sequence dimension. We then extract the features of the observed spectrum using a 3D transformer, with the query vector generated based on the representation of the molecular transition data. For the patch-level attention, the query vector is generated using only the embedding of the [CLS] token. In other words, the same query vector is shared across the sequence dimension. For the sequence-level attention, the query vector is generated per token but shared across the patch dimension. For the sequence of the molecular transition data, we employ two transformer layers (NIPS2017_3f5ee243) to model the correlation between the transitions. Unlike in NLP, our sequences are unordered, and therefore no positional embeddings are applied to the transformers. Finally, the [CLS] token embedding corresponding to the molecular spectroscopic data is passed to the decoder.
@@
The decoder of the neural network is a normalizing flow (e.g. NIPS2017_6c1da886; papamakarios2021normalizing), which models conditional probability distributions by iteratively applying reversible transformations to a simple multi-variate distribution. Normalizing flows allow both density estimation and sampling from the modeled distribution. This capability is essential for calculating the policy gradient during our training. In astronomy, normalizing flows have been applied to parameter estimation in cosmology (2019MNRAS.488.4440A) and gravitational wave signal analysis (2023PhRvD.108d2004B).
@@
This work employs the autoregressive flow, which is composed of a parameterized transformation and a conditioner (papamakarios2021normalizing). We adopt the piecewise quadratic transformation proposed by mueller18neural-v2 and the masked transformer (NIPS2017_3f5ee243) for the conditioner. Similar neural networks were presented by patacchiola2024transformer, who showed that the use of transformers could lead to better results in their benchmarks. When using the transformer, each input parameter is treated as a one-dimensional token, which is projected using a fully connected (FC) layer to match the dimension of the encoder embedding. Then, the embedding of each parameter is combined with the same encoder embedding via addition. For each token, the piecewise quadratic transformation is parameterized using $2N_{\text{bin}}+1$ dimensions, where $N_{\text{bin}}$ is the number of bins used to partition the domain of the piecewise polynomials. Consequently, we employ a FC layer to project the output of the masked transformer to the required $2N_{\text{bin}}+1$ dimensions. In this work, the decoder predicts five parameters, which are explained in Section 2.3.1. The output of the quadratic spline module is bounded, and the adopted bounds are given in Table 1.
@@
After several experiments, we suggest the following hyperparameters for our neural network. For the multi-head attention and transformer layers, we adopt 768 embedding dimensions and 8 heads. The hidden size in the MLP is set to be $768\times 8/3=2048$, a common practice when using the SwiGLU activation function (2020arXiv200205202S). For the sequence of the observed spectrum, the sub-windows are split into 8 patches. We adopt 32 bins for the quadratic spline module. In addition, as illustrated in Figure 1, we use 2 blocks and 1 block of the corresponding aggregated modules for the encoder and decoder, respectively. Our neural network has 89 million parameters in this configuration.
2.3 The spectral line fitting module
@@
2.3.1 The spectral line model
This work adopts the one-dimensional LTE spectral line model (see 2017A&A...598A...7M, for a nice derivation). In this work, we only consider the spectrum of a molecule or isotopologue at one vibrational state with one velocity component. In this case, the model spectrum is computed by

### v2

@@ Abstract
Spectral lines from interstellar molecules provide crucial insights into the physical and chemical conditions of the interstellar medium. Traditional spectral line analysis relies heavily on manual intervention, which becomes impractical when handling the massive datasets produced by modern facilities like ALMA. To address this challenge, we introduce a novel deep reinforcement learning framework to automate spectral line fitting. Using observational data from ALMA, we train a neural network that maps both molecular spectroscopic data and observed spectra to physical parameters such as excitation temperature and column density. The neural network predictions can serve as initial estimates and be further refined using a local optimizer. Our method achieves consistent fitting results compared to global optimization with multiple runs, while reducing the number of forward modeling runs by an order of magnitude. We apply our method to pixel-level fitting for an observation of the G327.3-0.6 hot core and validate our results using xclass. We perform the fitting for typical complex organic molecules of hot cores, including CH3OH, CH3OCHO, CH3OCH3, C2H5CN, and C2H3CN. For a 100 $\times$ 100 region covering 5 GHz bandwidth, the fitting process requires 4.9 to 41.9 minutes using a desktop with 16 cores and one consumer-grade GPU card.
@@
facilities: ALMA, SMA, IRAM:30msoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), casa (CASA Team et al., 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010; pandas development team, 2020), scipy (Virtanen et al., 2020), statcont (Sánchez-Monge et al., 2018), xclass (Möller et al., 2017).
I Introduction
@@
Spectral lines from interstellar species provide a crucial probe of the physical and chemical conditions in the interstellar medium (ISM). The analysis of spectral lines involves several key steps: querying transitions from spectroscopic databases, matching these transitions to observed spectral lines, and determining a set of physical parameters that best reproduce the observed spectrum (e.g. Goldsmith and Langer, 1999). Traditionally, spectral line analysis has been a labor-intensive process, requiring significant expertise and time to manually match observed features with known transitions and iteratively adjust parameters to achieve satisfactory fits. Although several analysis tools have been developed over the past decade, e.g. weeds (Maret et al., 2011), cassis (Vastel et al., 2015), xclass (Möller et al., 2017), madcuba (Martín et al., 2019), and pyspeckit (Ginsburg et al., 2022), manual intervention remains common. As the volume and complexity of observational data continue to increase due to advanced observational facilities like the Atacama Large Millimeter/submillimeter Array (ALMA) (Carpenter et al., 2020), there is a growing need for automated and efficient methods to perform this analysis, especially in line-rich regions such as hot cores and hot corinos (see Jørgensen et al., 2020, for a review).
@@
Recent advances in machine learning offer a promising opportunity to transform spectral analysis workflows. For instance, machine learning algorithms have been used to establish the relation between spectral line properties to the underlying physical parameters (Einig et al., 2024; Grassi et al., 2025; Mendoza et al., 2025). Kessler et al. (2025) trained a convolutional neural network for identification of complex organic molecules (COMs) using their spectral line features. In addition, machine learning techniques were applied to explore the chemical inventories of molecular clouds (Lee et al., 2021; Toru Shay et al., 2025) and to investigate the interstellar chemical reaction networks (Villadsen et al., 2022; Heyl et al., 2023).
@@
In this paper, we present a deep learning framework to automate and enhance the spectral line fitting process. Specifically, we aim to train a neural network that can infer the physical parameters of molecule emission such as excitation temperature and column density from the observed spectrum and the molecular transition properties within the observed spectral windows. We choose deep learning over traditional machine learning methods such as random forests (Breiman, 2001) because it scales more effectively with large datasets and is better suited to handling the massive data volumes produced by ALMA and other modern observatories. The inferred physical parameters of the neural network are based on a one-dimensional radiative transfer model that assumes local thermodynamic equilibrium (LTE). While this assumption may introduce bias (e.g. Roueff et al., 2024), the simplicity of the LTE model enables efficient computation of model spectra, which significantly reduces the time required for training and testing and makes the project feasible.
@@
This work employs deep reinforcement learning (DRL) (e.g. Mnih et al., 2015; Schulman et al., 2017) to train the neural network. To improve the generalization of the neural network, we seek to train it using real ALMA data. However, vast amounts of labeled data are unavailable for this task. This challenge could be addressed using DRL. Unlike supervised learning, DRL does not require labeled data and instead discovers optimal results during training. In our approach, training proceeds by interacting with a reward model and receiving feedback through reward signals that reflect fit quality. The reward model uses the physical parameters inferred by the neural network to compute the model spectra under LTE conditions and estimates the fit quality by comparing them to the observed spectrum.
@@
Our neural network-based approach can be efficiently applied to pixel-by-pixel fitting of spectral line cubes, which is a challenging task. When using local optimization algorithms, providing good initial parameter estimates for each pixel is non-trivial, and poor initial guesses can significantly affect the the fitting results. For instance, to improve the fitting, El-Abd et al. (2024) proposed using the results from previously fitted neighboring pixels to initialize the parameters for subsequent fits. On the other hand, while global optimization methods could lead to more robust fitting results, they are time consuming and require large number of forward modeling runs. In contrast, our neural network can provide good initial guesses for all pixels, leading to efficient and robust pixel-by-pixel fitting.
This paper is organized as follows. Section II introduces our methodology, including the description of the training data, along with the training and evaluation methods. Section III presents the training results. In Section IV, we apply our method to pixel-level spectral fitting. We then discuss our methodology in Section V. Finally, this work is summarized in Section VI.
@@
Figure 1: Architecture of the proposed neural network introduced in Section II.2. Here, FC stands for the fully connected layer, MLP stands for the multi-layer perceptron, and SwiGLU stands for the SwiGLU activation function (Shazeer, 2020).
II Methodology
II.1 Network inputs
@@
Given an observed spectrum with multiple spectral windows and a molecule identifier, we first query the molecular spectroscopic database to find all transitions of the molecule within those frequency ranges, resulting in a $N_{\text{trans}}\times F_{\text{trans}}$ array, denoted as $\textbf{d}_{\text{trans}}$, where $N_{\text{trans}}$ is the number of transitions and $F_{\text{trans}}$ is the number of features of each transition. For example, if the input spectral windows span 97 - 99 GHz and 100 - 102 GHz, and the target molecule is CH3OH, v=0, we query the molecular spectroscopic database for all transitions of CH3OH, v=0 within these frequency ranges, and may find 25 transitions; in this case $N_{\text{trans}}=25$. Data are grouped into batches during training. A single batch may contain multiple species, e.g. CO, v=0, CH3OH, v=0, and H2CO, v=0, and the corresponding spectral windows may also be different. Therefore, the number of transitions $N_{\text{trans}}$ varies across these input data. This situation is common in NLP, where the number of transitions is analogous to sentence length, and can be effectively handled using transformer-based neural networks (Vaswani et al., 2017). We adopt this network architecture in our work, as described in Section II.2.
We include the following features for each transition: the energy of the lower state $E_{\text{l}}$, the energy of the upper state $E_{\text{u}}$, the upper state degeneracy $g_{\text{u}}$, the Einstein coefficient $A_{\text{ul}}$, and the partition function Q at different temperatures. The molecular spectroscopic data are from the Virtual Atomic and Molecular Data Center (VAMDC) (Endres et al., 2016), which contains entries from the Cologne Database for Molecular Spectroscopy (CDMS) (Müller et al., 2001, 2005) and the Jet Propulsion Laboratory (JPL) database (Pickett et al., 1998).
@@
The third feature is derived using the approach described in Qiu et al. (2025). Specifically, we use the find_peaks function implemented by scipy, which identifies all local maxima in a 1D input array based on user-specified criteria. The most relevant criterion of the algorithm is the prominence threshold. Prominence is a property of a peak, defined as the vertical distance from the peak height to the lowest contour line that does not enclose any higher peak. The algorithm ignores any peak with a prominence below the specified threshold. In this work, we set the prominence threshold to four times the root mean square (RMS) noise of the input spectrum. The method to estimate the noise levels is described in Section II.4.1. In addition, the peak finding algorithm assumes that the input spectrum has a flat baseline. We have checked that this assumption is valid for the vast majority of the observational data used in this study.
@@
Table 1: Summary of the fitting parameters in our spectral line model described in Section II.3.1.
@@ II.2 Network architecture
Our network architecture, which consists of an encoder and a decoder, is illustrated in Figure 1. The neural network only uses basic modules in deep learning, including the fully connected (FC) layer, the multilayer perceptron (MLP), the patching layer in the vision transformer (ViT) (Dosovitskiy et al., 2021), the SwiGLU activation function (Shazeer, 2020), the multi-head attention module (Vaswani et al., 2017) and the transformer layer (Vaswani et al., 2017). Readers may also refer to Goodfellow et al. (2016) for the relevant concepts of deep learning.
@@
The encoder of our neural network converts the input variable-length sequences of molecular spectroscopic and observational data into dense vectors. This process is analogous to models in natural language processing (NLP) that learn vector embeddings of sentences and paragraphs, such as BERT (Devlin et al., 2019). A transformer-based architecture is particularly suitable for this purpose, and the so-called [CLS] token (Devlin et al., 2019) can be used to represent aggregate sequence-level information. In astronomy, transformer-based neural networks have been applied to learning the representations of galaxy images and spectra (e.g. Parker et al., 2024; Rizhko and Bloom, 2025).
@@
For the sequence of the observed spectrum, we first split the sub-window spectrum into several patches using the patching layer in ViT (Dosovitskiy et al., 2021). The same patching operation is applied across the sequence dimension. We then extract the features of the observed spectrum using a 3D transformer, with the query vector generated based on the representation of the molecular transition data. For the patch-level attention, the query vector is generated using only the embedding of the [CLS] token. In other words, the same query vector is shared across the sequence dimension. For the sequence-level attention, the query vector is generated per token but shared across the patch dimension. For the sequence of the molecular transition data, we employ two transformer layers (Vaswani et al., 2017) to model the correlation between the transitions. Unlike in NLP, our sequences are unordered, and therefore no positional embeddings are applied to the transformers. Finally, the [CLS] token embedding corresponding to the molecular spectroscopic data is passed to the decoder.
@@
The decoder of the neural network is a normalizing flow (e.g. Papamakarios et al., 2017, 2021), which models conditional probability distributions by iteratively applying reversible transformations to a simple multi-variate distribution. Normalizing flows allow both density estimation and sampling from the modeled distribution. This capability is essential for calculating the policy gradient during our training. In astronomy, normalizing flows have been applied to parameter estimation in cosmology (Alsing et al., 2019) and gravitational wave signal analysis (Bhardwaj et al., 2023).
@@
This work employs the autoregressive flow, which is composed of a parameterized transformation and a conditioner (Papamakarios et al., 2021). We adopt the piecewise quadratic transformation proposed by Müller et al. (2019) and the masked transformer (Vaswani et al., 2017) for the conditioner. Similar neural networks were presented by Patacchiola et al. (2024), who showed that the use of transformers could lead to better results in their benchmarks. When using the transformer, each input parameter is treated as a one-dimensional token, which is projected using a fully connected (FC) layer to match the dimension of the encoder embedding. Then, the embedding of each parameter is combined with the same encoder embedding via addition. For each token, the piecewise quadratic transformation is parameterized using $2N_{\text{bin}}+1$ dimensions, where $N_{\text{bin}}$ is the number of bins used to partition the domain of the piecewise polynomials. Consequently, we employ a FC layer to project the output of the masked transformer to the required $2N_{\text{bin}}+1$ dimensions. In this work, the decoder predicts five parameters, which are explained in Section II.3.1. The output of the quadratic spline module is bounded, and the adopted bounds are given in Table 1.
@@
After several experiments, we suggest the following hyperparameters for our neural network. For the multi-head attention and transformer layers, we adopt 768 embedding dimensions and 8 heads. The hidden size in the MLP is set to be $768\times 8/3=2048$, a common practice when using the SwiGLU activation function (Shazeer, 2020). For the sequence of the observed spectrum, the sub-windows are split into 8 patches. We adopt 32 bins for the quadratic spline module. In addition, as illustrated in Figure 1, we use 2 blocks and 1 block of the corresponding aggregated modules for the encoder and decoder, respectively. Our neural network has 89 million parameters in this configuration.
II.3 The spectral line fitting module
@@
II.3.1 The spectral line model
This work adopts a one-dimensional LTE spectral line model for a molecule or isotopologue in a single vibrational state with a single velocity component (see Möller et al., 2017, for a nice derivation). The model is characterized by five fitting parameters, the same quantities that our neural network predicts. Although real observed spectra typically contain emission from multiple species and vibrational states, fitting results for an individual species in a single vibrational state provide a valuable starting point for more detailed analysis. The resulting parameters could, for example, be used as initial values in subsequent joint fitting procedures.
The adopted spectral line model is computed by (Möller et al., 2017)